Circular No. 9085

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2009 U1 (GARRADD) G. J. Garradd, Siding Spring Observatory, reports his discovery of

a comet with the 0.5-m Uppsala Schmidt telescope (discovery observation tabulated below); stacking five 60-s exposures showed a 7′′.4 coma (FWHM) and faint fan tail 30′′ long in p.a. 160^{o} with a 45^{o} fan angle.

2009 UT α_{2000} δ_{2000} Mag. Observer
Oct. 17.71474 4^{h}25^{m}38.^{s}40 −39^{o}15′50′′.6 19.6 Garradd

Additional astrometry, the following very preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2009-U27:

T = 2010 July 7.819 TT  ω = 6.^{o}404
Ω = 67.196 }2000.0
q = 2.96357 AU  i = 68.944

COMET P/2009 U2 (LINEAR) J. V. Scotti reports his recovery of comet P/2001 YX_{127} (cf. IAUC

7828) with the Spacewatch II telescope at Kitt Peak; the comet was very slightly diffuse, with a short but faint tail. The indicated correction to the prediction in the 2009/2010 Comet Handbook is ΔT = −0.33 day (and ΔT = −0.36 day to the prediction on MPC 62881).

2009 UT α_{2000} δ_{2000} Mag. Observer
Oct. 18.45096 3^{h}23^{m}08.^{s}99 +21^{o}29′50′′.3 20.9 Scotti
18.46185 3 23 08.66 +21 29 50.2 21.0  _{”}
18.47364 3 23 08.29 +21 29 49.7 20.9  _{”}
19.27723 3 22 42.50 +21 29 33.1 21.2  _{”}
19.28834 3 22 42.12 +21 29 33.1 21.5  _{”}
19.30024 3 22 41.71 +21 29 32.4 21.0  _{”}

The following orbital elements (from observations 2001–2009), with residuals and an ephemeris, appear on MPEC 2009-U40.

Epoch = 2011 Aug. 27.0 TT

T = 2011 Aug. 23.8431 TT  ω = 114.^{o}7933
e = 0.177193  Ω = 31.0664 }2000.0
q = 3.430474 AU  i = 7.9154

a = 4.169232 AU n^{o} = 0.1157764 P = 8.51 years

2009 October 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green